# MCI-GRU: Stock Prediction Model Based on Multi-Head Cross-Attention and Improved GRU

### v1

@@ 4.2 Baseline Models
    Datasets CSI 300 CSI 500
    BLSW -0.076 0.113 -0.231 -0.670 -0.316 0.311 0.110 0.227 -0.155 0.485 0.710 0.446
    CSM -0.185 0.204 -0.293 -0.907 -0.631 -0.935 0.015 0.229 -0.179 0.066 0.084 0.001
    LSTM -0.214 0.175 -0.275 -1.361 -0.779 -1.492 -0.008 0.159 -0.172 -0.047 -0.044 -0.128
    ALSTM -0.216 0.164 -0.294 -1.314 -0.735 -1.461 0.016 0.162 -0.192 0.101 0.086 0.014
    GRU -0.229 0.156 -0.290 -1.469 -0.790 -1.631 -0.004 0.159 -0.193 -0.028 -0.023 -0.118
    Transformer -0.240 0.156 -0.281 -1.543 -0.855 -1.695 0.154 0.156 -0.135 0.986 1.143 0.867
    TRA -0.074 0.169 -0.222 -0.436 -0.332 -0.409 0.125 0.162 -0.145 0.776 0.866 0.657
    CTTS -0.193 0.206 -0.312 -0.937 -0.618 -0.907 -0.041 0.172 -0.239 -0.241 -0.173 -0.237
    A2C -0.207 0.092 -0.259 -2.255 -0.803 -2.490 -0.172 0.084 -0.208 -2.043 -0.826 -2.207
    DDPG -0.137 0.138 -0.240 -0.992 -0.568 -1.002 -0.128 0.082 -0.170 -1.563 -0.756 -1.639
    PPO -0.096 0.045 -0.120 -2.138 -0.800 -2.234 -0.032 0.015 -0.040 -2.041 -0.787 -2.075
    TD3 -0.154 0.137 -0.252 -1.122 -0.610 -1.155 -0.123 0.135 -0.248 -0.912 -0.496 -0.909
    SAC -0.140 0.090 -0.207 -1.554 -0.676 -1.635 -0.167 0.081 -0.207 -2.057 -0.807 -2.219
    FactorVAE -0.048 0.134 -0.175 -0.335 -0.271 -0.348 0.006 0.127 -0.147 0.047 0.041 0.112
    AlphaStock -0.164 0.153 -0.245 -1.072 -0.669 -1.098 -0.017 0.148 -0.166 0.115 -0.102 -0.043
    DeepPocket -0.036 0.135 -0.175 -0.270 -0.207 -0.258 0.006 0.127 -0.148 0.050 0.043 0.115
    DeepTrader -0.122 0.147 -0.229 -0.828 -0.533 -0.876 0.055 0.168 -0.141 0.324 0.388 0.370
    THGNN -0.015 0.172 -0.152 -0.088 -0.100 -0.003 0.048 0.128 -0.141 0.375 0.340 0.432
@@
BLSW 0.199 0.318 -0.223 0.626 0.892 0.774 0.368 0.339 -0.222 1.086 1.658 1.194
CSM 0.099 0.250 -0.139 0.396 0.712 0.584 0.116 0.242 -0.145 0.479 0.800 0.603
LSTM 0.142 0.162 -0.178 0.877 0.798 0.929 0.247 0.176 -0.128 1.403 1.930 1.386
ALSTM 0.191 0.161 -0.150 1.186 1.273 1.115 0.201 0.192 -0.183 1.047 1.098 1.032
GRU 0.124 0.169 -0.139 0.734 0.829 1.023 0.225 0.188 -0.165 1.197 1.364 1.160
Transformer 0.135 0.159 -0.140 0.852 0.968 0.908 0.268 0.175 -0.131 1.531 2.046 1.441
TRA 0.184 0.166 -0.158 1.114 1.172 1.106 0.267 0.181 -0.144 1.475 1.854 1.427
CTTS 0.154 0.161 -0.113 0.952 1.356 0.965 0.349 0.197 -0.193 1.769 1.808 1.610
A2C 0.160 0.126 -0.084 1.267 1.907 1.244 0.109 0.134 -0.114 0.816 0.957 0.844
DDPG 0.111 0.129 -0.091 0.864 1.223 0.887 0.130 0.156 -0.131 0.832 0.994 0.863
PPO 0.020 0.089 -0.067 0.220 0.291 0.263 0.148 0.118 -0.104 1.259 1.424 1.237
TD3 0.024 0.113 -0.105 0.209 0.225 0.264 0.181 0.155 -0.160 1.169 1.130 1.156
SAC 0.140 0.111 -0.069 1.263 2.011 1.242 0.162 0.139 -0.107 1.165 1.518 1.154
FactorVAE 0.160 0.142 -0.132 1.128 1.211 1.013 0.356 0.159 -0.119 2.234 2.995 1.907
AlphaStock 0.122 0.140 -0.126 0.871 0.968 0.892 0.372 0.178 -0.134 1.781 2.776 1.869
DeepPocket 0.165 0.142 -0.126 1.165 1.311 1.045 0.346 0.157 -0.116 2.197 2.971 1.882
DeepTrader 0.295 0.180 -0.181 1.635 1.628 1.425 0.716 0.248 -0.138 2.890 5.169 2.306
THGNN 0.271 0.141 -0.094 1.921 2.871 1.778 0.644 0.204 -0.146 3.147 3.414 2.543
@@ 4.3 Experimental Results
In this section, we conduct a rigorous evaluation of the experimental results of our proposed model in comparison with several baseline models across different datasets, as illustrated in Tables 2 and 3.
@@
On the CSI 300 dataset, traditional and deep learning models, including BLSW, CSM, LSTM, ALSTM, GRU, and Transformer, generally demonstrate subpar performance, characterized by negative ARR values and low ASR, CR, and IR metrics. For instance, while the Transformer model achieves the highest ARR among the traditional models, it still reports a negative ARR of -0.240 with a maximum drawdown (MDD) of -0.281, indicating considerable risk exposure. In the case of the CSI 500 dataset, performance improves slightly, with the Transformer model recording a positive ARR of 0.154 and an ASR of 0.986. However, these values remain notably lower than those delivered by our proposed model. Although reinforcement learning models, such as TRA, CTTS, A2C, DDPG, PPO, TD3, SAC, and FactorVAE, show marginal improvements, they continue to underperform relative to our model. For example, the PPO model reports the lowest annualized volatility (AVoL) and MDD, signaling lower risk, yet its ARR and IR metrics remain negative across both datasets, underscoring its lackluster return potential. Graph-based models, including AlphaStock, DeepPocket, DeepTrader, and THGNN, show more promising results, with THGNN, for instance, achieving an ARR of -0.015 on CSI 300 and 0.048 on CSI 500. Despite these improvements, their performance still falls short of our model.
@@ 4.5 Ablation Study
In this section, we conduct a comprehensive ablation study to evaluate the individual contributions of various components of our model. The model is systematically divided into four distinct modules: Use Improved GRU to Capture Temporal Features (I), Use GAT to Capture Cross-sectional Features (II), Use Multi-head Cross-attention to Capture Latent State Features (III), Model Prediction and Loss Calculation Layer (IV). Detailed experimental results are presented in Tables 10 and 11.
@@
The results presented in Tables 10 and 11 reveal several critical insights. First, the integration of the I and II resulted in a moderate improvement in performance metrics across all datasets, indicating that these components play a significant role in capturing both temporal and relational dependencies within the data. However, the addition of the III further amplified the model’s performance, suggesting that the incorporation of market-wide latent states provides a more comprehensive understanding of broader market dynamics. This enhancement underscores the importance of integrating diverse levels of information to effectively model complex financial relationships.

### v2

@@ 4.2 Baseline Models
  • 19.
    VGNN [89]: A decoupled graph learning framework for vague graph learning in stock prediction, leveraging matrix/tensor fusion, hybrid attention, and message passing to outperform state-of-the-art methods.
  • 20.
    MASTER [63]: A market-centric transformer model that dynamically simulates both instantaneous and cross-temporal correlations between stocks, enhancing trend forecast precision.
  • 21.
    CISTHPAN [90]: A two-stage framework combining Transformer and HGAT-based pre-training with stock-ranking fine-tuning, designed to improve the accuracy of stock movement predictions.
  • 22.
    CL [91]: A model combining textual and quantitative indicators for improved explainability in stock movement prediction, achieving state-of-the-art performance.
Datasets CSI 300 CSI 500
Model ARR $\uparrow$ AVol $\downarrow$ MDD $\downarrow$ ASR $\uparrow$ CR $\uparrow$ IR $\uparrow$ MSE $\downarrow$ MAE $\downarrow$ ARR $\uparrow$ AVol $\downarrow$ MDD $\downarrow$ ASR $\uparrow$ CR $\uparrow$ IR $\uparrow$ MSE $\downarrow$ MAE $\downarrow$
BLSW -0.076 0.113 -0.231 -0.670 -0.316 0.311 0.078 0.090 0.110 0.227 -0.155 0.485 0.710 0.446 0.033 0.064
CSM -0.185 0.204 -0.293 -0.907 -0.631 -0.935 0.093 0.124 0.015 0.229 -0.179 0.066 0.084 0.001 0.049 0.098
LSTM -0.214 0.175 -0.275 -1.361 -0.779 -1.492 0.108 0.142 -0.008 0.159 -0.172 -0.047 -0.044 -0.128 0.065 0.118
ALSTM -0.216 0.164 -0.294 -1.314 -0.735 -1.461 0.112 0.148 0.016 0.162 -0.192 0.101 0.086 0.014 0.045 0.093
GRU -0.229 0.156 -0.290 -1.469 -0.790 -1.631 0.123 0.155 -0.004 0.159 -0.193 -0.028 -0.023 -0.118 0.062 0.112
Transformer -0.240 0.156 -0.281 -1.543 -0.855 -1.695 0.131 0.159 0.154 0.156 -0.135 0.986 1.143 0.867 0.025 0.053
TRA -0.074 0.169 -0.222 -0.436 -0.332 -0.409 0.079 0.088 0.125 0.162 -0.145 0.776 0.866 0.657 0.030 0.062
CTTS -0.193 0.206 -0.312 -0.937 -0.618 -0.907 0.095 0.129 -0.041 0.172 -0.239 -0.241 -0.173 -0.237 0.078 0.131
A2C -0.207 0.092 -0.259 -2.255 -0.803 -2.490 0.099 0.137 -0.172 0.084 -0.208 -2.043 -0.826 -2.207 0.098 0.166
DDPG -0.137 0.138 -0.240 -0.992 -0.568 -1.002 0.087 0.106 -0.128 0.082 -0.170 -1.563 -0.756 -1.639 0.088 0.145
PPO -0.096 0.045 -0.120 -2.138 -0.800 -2.234 0.083 0.096 -0.032 0.015 -0.040 -2.041 -0.787 -2.075 0.073 0.127
TD3 -0.154 0.137 -0.252 -1.122 -0.610 -1.155 0.090 0.115 -0.123 0.135 -0.248 -0.912 -0.496 -0.909 0.085 0.139
SAC -0.140 0.090 -0.207 -1.554 -0.676 -1.635 0.086 0.109 -0.167 0.081 -0.207 -2.057 -0.807 -2.219 0.095 0.158
FactorVAE -0.048 0.134 -0.175 -0.335 -0.271 -0.348 0.071 0.085 0.006 0.127 -0.147 0.047 0.041 0.112 0.055 0.103
AlphaStock -0.164 0.153 -0.245 -1.072 -0.669 -1.098 0.092 0.112 -0.017 0.148 -0.166 0.115 -0.102 -0.043 0.069 0.122
DeepPocket -0.036 0.135 -0.175 -0.270 -0.207 -0.258 0.065 0.080 0.006 0.127 -0.148 0.050 0.043 0.115 0.056 0.104
DeepTrader -0.122 0.147 -0.229 -0.828 -0.533 -0.876 0.084 0.102 0.055 0.168 -0.141 0.324 0.388 0.370 0.039 0.072
THGNN -0.015 0.172 -0.152 -0.088 -0.100 -0.003 0.043 0.071 0.048 0.128 -0.141 0.375 0.340 0.432 0.045 0.078
VGNN -0.037 0.163 -0.197 -0.227 -0.188 -0.201 0.061 0.082 0.111 0.166 -0.175 0.668 0.636 0.564 0.031 0.063
MASTER 0.102 0.151 -0.126 0.681 0.811 0.726 0.057 0.078 0.128 0.130 -0.098 0.989 1.308 0.997 0.028 0.057
CI-STHPAN -0.078 0.167 -0.144 -0.466 -0.538 -0.355 0.075 0.092 0.021 0.151 -0.129 0.136 0.161 0.211 0.053 0.088
CL -0.035 0.148 -0.183 -0.241 -0.195 -0.193 0.063 0.078 0.051 0.146 -0.128 0.351 0.401 0.390 0.041 0.075
MCI-GRU 0.352 0.226 -0.127 1.559 2.776 1.526 0.035 0.068 0.330 0.203 -0.198 1.626 1.663 1.382 0.022 0.051
@@
Model ARR $\uparrow$ AVol $\downarrow$ MDD $\downarrow$ ASR $\uparrow$ CR $\uparrow$ IR $\uparrow$ MSE $\downarrow$ MAE $\downarrow$ ARR $\uparrow$ AVol $\downarrow$ MDD $\downarrow$ ASR $\uparrow$ CR $\uparrow$ IR $\uparrow$ MSE $\downarrow$ MAE $\downarrow$
BLSW 0.199 0.318 -0.223 0.626 0.892 0.774 0.068 0.094 0.368 0.339 -0.222 1.086 1.658 1.194 0.058 0.082
CSM 0.099 0.250 -0.139 0.396 0.712 0.584 0.125 0.173 0.116 0.242 -0.145 0.479 0.800 0.603 0.119 0.178
LSTM 0.142 0.162 -0.178 0.877 0.798 0.929 0.097 0.132 0.247 0.176 -0.128 1.403 1.930 1.386 0.085 0.129
ALSTM 0.191 0.161 -0.150 1.186 1.273 1.115 0.073 0.098 0.201 0.192 -0.183 1.047 1.098 1.032 0.084 0.147
GRU 0.124 0.169 -0.139 0.734 0.829 1.023 0.109 0.151 0.225 0.188 -0.165 1.197 1.364 1.160 0.089 0.138
Transformer 0.135 0.159 -0.140 0.852 0.968 0.908 0.103 0.145 0.268 0.175 -0.131 1.531 2.046 1.441 0.082 0.121
TRA 0.184 0.166 -0.158 1.114 1.172 1.106 0.079 0.104 0.267 0.181 -0.144 1.475 1.854 1.427 0.081 0.119
CTTS 0.154 0.161 -0.113 0.952 1.356 0.965 0.093 0.125 0.349 0.197 -0.193 1.769 1.808 1.610 0.071 0.102
A2C 0.160 0.126 -0.084 1.267 1.907 1.244 0.089 0.117 0.109 0.134 -0.114 0.816 0.957 0.844 0.126 0.182
DDPG 0.111 0.129 -0.091 0.864 1.223 0.887 0.121 0.165 0.130 0.156 -0.131 0.832 0.994 0.863 0.110 0.176
PPO 0.020 0.089 -0.067 0.220 0.291 0.263 0.159 0.212 0.148 0.118 -0.104 1.259 1.424 1.237 0.104 0.165
TD3 0.024 0.113 -0.105 0.209 0.225 0.264 0.151 0.201 0.181 0.155 -0.160 1.169 1.130 1.156 0.093 0.160
SAC 0.140 0.111 -0.069 1.263 2.011 1.242 0.099 0.136 0.162 0.139 -0.107 1.165 1.518 1.154 0.098 0.158
FactorVAE 0.160 0.142 -0.132 1.128 1.211 1.013 0.088 0.115 0.356 0.159 -0.119 2.234 2.995 1.907 0.063 0.089
AlphaStock 0.122 0.140 -0.126 0.871 0.968 0.892 0.112 0.157 0.372 0.178 -0.134 1.781 2.776 1.869 0.053 0.084
DeepPocket 0.165 0.142 -0.126 1.165 1.311 1.045 0.084 0.112 0.346 0.157 -0.116 2.197 2.971 1.882 0.061 0.098
DeepTrader 0.295 0.180 -0.181 1.635 1.628 1.425 0.049 0.082 0.716 0.248 -0.138 2.890 5.169 2.306 0.019 0.048
THGNN 0.271 0.141 -0.094 1.921 2.871 1.778 0.057 0.086 0.644 0.204 -0.146 3.147 3.414 2.543 0.026 0.057
VGNN 0.299 0.202 -0.169 1.473 1.767 1.406 0.045 0.076 0.616 0.181 -0.091 3.405 6.767 2.798 0.033 0.063
MASTER 0.335 0.171 -0.134 1.958 2.495 1.895 0.038 0.065 0.654 0.168 -0.076 3.876 8.534 3.183 0.023 0.052
CI-STHPAN 0.123 0.233 -0.254 0.527 0.485 0.632 0.114 0.159 0.454 0.208 -0.124 2.178 3.660 1.855 0.041 0.071
CL 0.308 0.189 -0.171 1.629 1.794 1.451 0.039 0.071 0.351 0.172 -0.122 2.041 2.877 1.821 0.066 0.093
MCI-GRU 0.456 0.179 -0.129 2.549 3.543 2.197 0.031 0.062 0.718 0.220 -0.118 3.257 6.091 2.609 0.015 0.039
@@ 4.3 Experimental Results
In this section, we conduct a rigorous evaluation of the experimental results of our proposed model in comparison with several baseline models across different datasets, as illustrated in Tables 3 and 4.
@@
On the CSI 300 dataset, traditional and deep learning models (including BLSW, CSM, LSTM, ALSTM, GRU, and Transformer) generally exhibit subpar performance, characterized by negative ARR values and low ASR, CR, and IR metrics. For example, although the Transformer model achieves the highest ARR among the traditional models, its ARR is still -0.240, with an MDD of -0.281, indicating significant risk exposure. In the case of the CSI 500 dataset, the performance shows slight improvement, with the Transformer model achieving an ARR of 0.154 and an ASR of 0.986. However, these values are still significantly lower than those offered by our proposed model. While reinforcement learning models (such as TRA, CTTS, A2C, DDPG, PPO, TD3, SAC, and FactorVAE) show marginal improvements, their performance remains below that of our model. For instance, the PPO model has the lowest AVoL and MDD, indicating lower risk, but its ARR and IR metrics are still negative in both datasets, highlighting poor return potential. Graph-based models (including AlphaStock, DeepPocket, DeepTrader, THGNN, VGNN, MASTER, and CISTHPAN) show more promising results. For example, THGNN achieves an ARR of -0.015 on the CSI 300 and an ARR of 0.048 on the CSI 500. Among these graph-based models, VGNN demonstrates relatively good ARR values on both datasets, with -0.037 on the CSI 300 and 0.111 on the CSI 500, but its risk-adjusted return metrics (ASR, CR, IR) still require improvement. The MASTER model achieves a relatively high ARR of 0.102 on the CSI 300, along with relatively high ASR, CR, and IR, showing some potential, but its performance on the CSI 500 is comparatively mediocre. The CI-STHPAN model performs poorly on both datasets, with low ARR and high-risk metrics. The CL model performs poorly on the CSI 300 but shows some improvement on the CSI 500 with an ARR of 0.051, though its volatility is also relatively high. Despite their individual characteristics and performances, the performance of these graph-based models still falls short compared to our model.
Overall, the model performs better in the U.S. market than in the Chinese market based on the ARR metric. In particular, the NASDAQ 100 index shows high ARR values, with MCI-GRU reaching 0.718, reflecting strong growth in the U.S. tech market. In contrast, ARR values in the Chinese market are generally lower, and many traditional models yield negative ARR for the CSI 300 index. This discrepancy may stem from higher volatility, a more complex market structure, and investor behavior in China. Even the best-performing MCI-GRU achieves only 0.352 on the CSI 300 index, lower than in the U.S. The U.S. market exhibits relatively low volatility, particularly in the NASDAQ 100, where many models have aAVol below 0.2, indicating stability. Conversely, the Chinese market shows higher volatility, likely due to its younger market environment and dominance of retail investors, which complicates prediction and risk management. Regarding MDD, the U.S. market typically shows smaller MDD values, reinforcing its stability, while the Chinese market experiences higher MDD, indicating greater downside risk. In terms of risk-adjusted metrics—ASR, CR, and IR—the U.S. market shows higher values, driven by higher ARR and lower volatility, suggesting better risk-adjusted returns. The Chinese market has lower risk-adjusted returns, primarily due to lower ARR and higher volatility. The MCI-GRU model performs well across all four datasets, demonstrating robustness and adaptability in different market environments. In contrast, traditional and reinforcement learning models perform notably worse in the Chinese market, indicating they may be better suited for more mature, stable markets.
Additionally, as shown in Tables 3 and 4, there are significant differences in the performance of various models in terms of Mean Squared Error (MSE) and Mean Absolute Error (MAE). Since this study focuses primarily on predicting daily stock returns, and the prediction errors are relatively small, the resulting MSE and MAE values are correspondingly modest. Although MAE and MSE are commonly used metrics for evaluating prediction accuracy, there is some correlation between these metrics and investment return indicators (such as ARR, ASR, and IR). In general, models with higher accuracy are likely to yield better investment returns. However, it is important to note that lower MAE and MSE do not necessarily translate into higher investment returns, as returns are influenced by various factors, including trading strategies and risk control. Among all the models, MCI-GRU performs exceptionally well in terms of MAE and MSE, particularly on the NASDAQ 100 dataset, where its MAE and MSE are the lowest, indicating a clear advantage in prediction accuracy. Its excellent performance in investment return metrics further supports the correlation between higher prediction accuracy and better investment returns. In contrast, some traditional and reinforcement learning models perform relatively poorly in terms of MAE and MSE, which is consistent with their performance in investment return metrics, suggesting that these models may not effectively capture the dynamic fluctuations of the market. GNN-based models generally show better performance in terms of MAE and MSE, indicating that graph structural information helps improve prediction accuracy. Among these models, DeepTrader, THGNN, VGNN, and MASTER perform well on both MAE and MSE, further demonstrating the importance of graph structures in capturing complex relationships in the stock market and enhancing prediction accuracy.
@@
The model integrates several complex components, including the improved GRU, GAT, and multi-head cross-attention mechanisms, which may lead to high computational costs. Specifically, the modified GRU replaces the reset gate in the traditional GRU, and its computational complexity is $O(d\cdot n)$, where $d$ is the hidden state dimension and $n$ is the number of time steps. The computational complexity of the Graph Attention Network (GAT) is $O(|E|\cdot d)$, where $|E|$ is the number of edges in the graph and $d$ is the feature dimension of each node. Since GAT requires calculating attention weights between nodes, its computational complexity is closely related to the sparsity of the graph and the number of nodes. Additionally, the multi-head cross-attention mechanism introduces a computational complexity of $O(n\cdot d_{i}^{2})$ for each attention head, with a total complexity of $O(h\cdot n\cdot d_{i}^{2})$, where $h$ is the number of attention heads and $d_{i}$ is the dimension of each head. When the number of attention heads $h$ or the dimension $d_{i}$ is large, the computational cost can increase significantly. To address this challenge and improve the model’s computational efficiency, several optimization strategies can be employed. For instance, parameter sharing and low-rank approximation techniques can be used in both the GRU and multi-head cross-attention mechanisms to reduce the computational burden. Additionally, GAT’s graph structure can be optimized by exploiting sparsity, reducing unnecessary edges, or applying neighborhood sampling to lower the computational complexity. The multi-head attention mechanism can be accelerated by parallelizing the computation of attention heads, reducing overall computation time. Although the model has a high computational cost, its complexity can be effectively controlled through these optimization strategies.
@@ 4.5 Ablation Study
In this section, we conduct a comprehensive ablation study to evaluate the individual contributions of various components of our model. The model is systematically divided into four distinct modules: Use Improved GRU to Capture Temporal Features (I), Use GAT to Capture Cross-sectional Features (II), Use Multi-head Cross-attention to Capture Latent State Features (III), Model Prediction and Loss Calculation Layer (IV). Detailed experimental results are presented in Tables 11 and 12.
@@
The results presented in Tables 11 and 12 reveal several critical insights. First, the integration of the I and II resulted in a moderate improvement in performance metrics across all datasets, indicating that these components play a significant role in capturing both temporal and relational dependencies within the data. However, the addition of the III further amplified the model’s performance, suggesting that the incorporation of market-wide latent states provides a more comprehensive understanding of broader market dynamics. This enhancement underscores the importance of integrating diverse levels of information to effectively model complex financial relationships.
@@ 4.6 Case studies
4.7 Limitations of the model
Although the proposed MCI-GRU model has demonstrated excellent performance in multiple experiments and has been successfully applied in a fund management company, we have not discussed its potential limitations in detail. In terms of scalability, as the dataset size increases, the training time and computational resource consumption of the MCI-GRU model may grow, especially when dealing with a large number of stocks or multiple markets. While we have achieved good results on the Chinese and U.S. stock market datasets, the training and inference efficiency of the model may be impacted as the data scale expands. Future work will focus on improving scalability through methods such as optimizing computational graphs, model compression, or distributed training.
Regarding robustness to market volatility, although the model performs well on historical data, its robustness under extreme market fluctuations still needs further validation. In real-world applications, severe market volatility may lead to model prediction failure. Therefore, we plan to explore the model’s adaptability in different market environments and consider incorporating risk management mechanisms to enhance the model’s robustness to market volatility.
As for sensitivity to hyperparameter tuning, although we have selected appropriate hyperparameters in our experiments and validated the model’s robustness to some extent, MCI-GRU may still be highly sensitive to certain hyperparameters (such as learning rate, number of attention heads, etc.). To reduce this sensitivity, future research will explore adaptive hyperparameter tuning methods or use techniques like Bayesian optimization to further improve the model’s generalization ability.